OpenStudy (anonymous):

T(My-Nx)=NTx-MTy find T? T(x,y),M(x,y) and N(x,y). My represents here partial derivative of M w.r.t y, and similarly others.

myininaya (myininaya):

\[T(M_y-N_x)=NT_x-MT_y\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any idea??

OpenStudy (anonymous):

I tried to solve by T=X(x) Y(y) but did'nt work

myininaya (myininaya):

oh i remember that way thats my favorite way because its easiest i was just looking in my partial differential equation book because i do not remember how

OpenStudy (anonymous):

OK

myininaya (myininaya):

but what did you do with N and M?

OpenStudy (anonymous):

I leave them. assume them as function of x and y

myininaya (myininaya):

help me think i don't know if i'm going anywhere but i did this: Let T=X(x)Y(x), M=G(x)F(y), N=H(x)L(y) ... then I eventually got XG(YF)'=LY(HX)' but I don't know if this is gonna take me anywhere

OpenStudy (anonymous):

you are assuming T as a function of x, in that case T=exp(-MTy/My-Nx)

myininaya (myininaya):

we can put all of our functions of y on one side and all of our functions of x on the other side (YF)'/(LY)=(HX)'/(XG) ...

OpenStudy (anonymous):

M and N will be given so we don't need to decompose them

OpenStudy (anonymous):

we just want to find out T

OpenStudy (anonymous):

And I want to solve for case when T(x,y)

myininaya (myininaya):

jamal i give up for now i'm sorry

OpenStudy (anonymous):

I found one T=yf(x) then T=yexp((y(My-Nx)+M)/N) but this is not the most genral case

OpenStudy (anonymous):

And I found this solution by discussing you thank you.

myininaya (myininaya):

i need to review my differential/partial differential equations :(

myininaya (myininaya):

you're so sweet jamal that makes me feel better

myininaya (myininaya):

:)